# Unraveling the Black-box Magic: An Analysis of Neural Networks’ Dynamic Extrema

### v2

@@ Abstract
We point out that neural networks are not black boxes, and their generalization stems from the ability to dynamically map a dataset to the extrema of the model function. We further prove that the number of extrema in a neural network is positively correlated with the number of its parameters. We then propose a new algorithm that is significantly different from back-propagation algorithm, which mainly obtains the values of parameters by solving a system of linear equations. Some difficult situations, such as gradient vanishing and overfitting, can be reasonably explained and dealt with in this framework.
Keywords: neural network, generalization, black box, extreme increment, linear equation system
@@ 2.4 General Training Process of an Ideal Model
The model trained through the above steps can accurately map the input to the output, thereby enabling the machine to precisely answer “yes” or “no”. If in actual debugging, we cannot find or it is very difficult to find an ideal function like that in Figure 5, then a weakened function is trained that the surfaces or points within the surfaces’ neighbors is adjusted to be its local extreme points first, and then the parameters of the function are adjusted to move these local extreme points as close as possible to the upper or lower limit of the function’s range, thereby achieving a function that is close to that in Figure 5.
@@ 4.3 Reduce the Computational Scale
4) After the training is completed, verify whether the predicted values of all non-central samples on the neural network are within the specified accuracy range. If the accuracy requirements are met, the algorithm ends. Otherwise, mark the non-central samples that do not meet the requirements as central samples and repeat steps 3) and 4).
@@ 6.1 Polarization
Given a set of binary classification functions $\{h_{v}^{[n]}(x)|v\in[1,l_{n}]\}$, how to flip and expand their local extreme points?
For example, for the parabolic function $y=a{(x-b)}^{2}$, by changing the sign of the parameter $a$, we can flip the extreme point, and by increasing the value of parameter $a$, the extreme value will increase. Furthermore, apart from $L(n,-,\Phi)$, $L(u,-,\Phi)$ with $u\in[1,n-1]$ are all atypical homogeneous linear equations. Whether their unique structure is conducive to obtaining a particular solution is also worthy of further discussion.

### v3

# Unraveling the Black Box of Neural Networks: A Dynamic Extremum Mapper

@@ Abstract
We point out that neural networks are not black boxes, and their generalization stems from the ability to dynamically map a dataset to the extrema of the model function. We further prove that the number of extrema in a neural network is positively correlated with the number of its parameters. We then propose a new algorithm that is significantly different from back-propagation algorithm, which mainly obtains the values of parameters by solving a system of linear equations. Some difficult situations, such as gradient vanishing and overfitting, can be simply explained and dealt with in this framework.
Keywords: neural network, generalization, black box, extreme increment, homogeneous system of linear equations, large language model
@@ 2.4 General Training Process of an Ideal Model
The model trained through the above steps can accurately map the input to the output, thereby enabling the machine to precisely answer “yes” or “no”. In Section 3, it was demonstrated that neural networks can be decomposed into a set of binary classification functions, with the corresponding surfaces mapped to the local extrema of these functions by finding the general solution of a homogeneous system of linear equations, thereby establishing the validity of Steps 1 and 2. In Section 4, a method for mapping the surfaces to the global extrema of the functions by enumerating the particular solutions of the homogeneous system of linear equations was presented, thus confirming Steps 3 and 4.
@@ 4.3 Reduce the Computational Scale
4) After the training is completed, verify whether the predicted values of all non-central samples on the neural network are within the specified accuracy range. If the accuracy requirements are met, the algorithm ends. Otherwise, mark the non-central samples that do not meet the requirements as central samples and repeat steps 3 and 4.
@@ 6.1 Polarization
Given a set of binary classification functions $\{h_{v}^{[n]}(x)|v\in[1,l_{n}]\}$, how to flip their local extreme points and expand the values arbitrarily?
For example, for a parabolic function $y=a{(x-b)}^{2}$, by changing the sign of the parameter $a$, we can flip its extreme point, and by increasing the value of parameter $a$, the extreme value will increase. Furthermore, apart from $L(n,-,\Phi)$, $L(u,-,\Phi)$ with $u\in[1,n-1]$ are all atypical homogeneous linear equations. Whether their unique structure is conducive to obtaining a particular solution is also worthy of further discussion.
@@ 7 Summary
Acknowledgments and Disclosure of Funding
I am extremely grateful to my supervisor, Professor Yang Lihua, who provided me with the initial guidance and assistance on the basic research direction when I first started my research journey. Without his guidance, I might have chosen a different research path, thus missing out on this study. Additionally, this research was funded by the Science and Technology Research Project of the Key Areas in Nanhai District, Foshan City (Grant No. 2230032004637). I would like to express my gratitude for this as well.